OpenStudy (anonymous):

find the partial derivatives with respect to x and y, if f(x,y) = \[\sum_{n=0}^{}\] (xy)^n

OpenStudy (anonymous):

|xy| <1

OpenStudy (anonymous):

So, wrt to one variable, you treat the other as a constant. So how would you differentiate (constant*x)^n?

OpenStudy (anonymous):

(n)x^(n-1)? but what about the sum?

OpenStudy (anonymous):

@First question - Yep. @Second question - What's the upper limit? Also, what's the |xy| < 1 for? Sorry, I'm not quite understanding the question fully.

OpenStudy (anonymous):

the upper limit is infinity. |xy|<1 is the condition that goes along with the question

OpenStudy (anonymous):

Ah! Makes sense. So, basically, for x, it would be: \[ \displaystyle\sum_{n=0}^{\infty} nx^{n-1}. \]At this point, I'll just wait for another response. I really have no idea myself. Sorry. :( All the best!

OpenStudy (anonymous):

patial derivative wrt to x \[\sum_{n=0}^{n=\infty}nx ^{n-1}y ^{n}\]

OpenStudy (anonymous):

and wrt to y \[\sum_{n=0}^{n=\infty}x ^{n}(ny ^{n-1})\]

OpenStudy (anonymous):

any query?

OpenStudy (anonymous):

I dont think that is quite right. I dont get how you take a partial of a sum and how you would account for |xy| < 1

OpenStudy (anonymous):

ok first tell me the partial derivative of xy^3

OpenStudy (anonymous):

wrt to x??????

OpenStudy (anonymous):

just find out and tell me

OpenStudy (anonymous):

find 2 different partials, one with relation to x and one with relation to y

OpenStudy (anonymous):

ya u find